The following results came from a simple regression analysis based on 20 oberservations.

Yhat=15-5x

SSyy=400

Error Sum of square=100

Source SS df MS F
regression 300.00 1 300.00 54.00
error 100.00 18 5.56
total 400.00 19

What is the numerical value for the sample correlation coefficient, and how many standard deviations away from zero is the value you used as your previous answer? .
Answer #1

Coefficient of Determination = r^2  
= SS(Regression) / SS(Total)
= 300.00/ 400.00 = 3/4 = 0.75

correlation coefficient = r = 0.8660

Standard deviation is = √SS(error)/n-2
= 2.3570

So there are 2 standard deviations is away from zero